# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ IV-E Poisoning & Backdoor Attacks in LM Agent Networks
Different from adversarial attacks, poisoning and backdoor attacks typically involve altering model parameters by injecting toxic data into the training dataset during model training process, thereby deteriorating model performance or injecting specific backdoors. In the following, we review the latest advances on poisoning attack and backdoor attacks to LM agents.
IV-E1 Poisoning Attacks
@@
Poisoning attacks refer to the manipulation of a model’s behavior by introducing toxic information, such as malicious training data, resulting in the degradation of model’s generalization ability or generation of predetermined errors for specific inputs. For LM agents, poisoning attacks include both conventional forms (e.g., data poisoning and model poisoning) and new techniques tailored to LM agents (e.g., RAG poisoning and agent poisoning).
@@
  • Data poisoning: Data poisoning is the most common form of poisoning attacks. The risk of data poisoning is significantly increased in LM agents due to the presence of a large amount of unverified data collected from the Internet and user-model interactions. Scheuster et al. [164] show that neural code autocompleters are vulnerable to poisoning attacks. By introducing toxic files to the autocompleter’s training corpus, adversaries can manipulate the autocompleter to generate insecure suggestions. Wan et al. [165] demonstrate that during the instruction tuning process, adversaries can exploit poisoned examples featuring a specific trigger phrase to induce frequent misclassifications or degrade the quality of outputs.
@@
  • Model poisoning: As depicted in Fig. 25, in distributed model pre-tuning/inference paradigms such as FL for cooperative LM agents, attackers can impersonate benign agents and upload poisoned model updates during each round of communication, thereby deteriorating the performance of the global LM [37]. For LM agents, leveraging cloud-edge collaboration for LM fine-tuning emerges as an effective way to harness the resources of distributed edge nodes [85]. However, the presence of malicious edge LM agents may pose risks of model poisoning during the collaborative training process, thereby compromising the global LM performance.
  • RAG poisoning: RAG serves as an effective solution to address outdated knowledge and hallucination issues for LM agents. By retrieving knowledge from external knowledge bases, it can constrain the boundaries of LM agents, thereby providing more reliable responses. However, as depicted in Fig. 26, attackers can perform poisoning attacks on knowledge bases, causing LM agents to generate unexpected responses. For instance, during code generation, the check for divisor zero may be omitted because the poisoned base contains the knowledge that 0 is correct as a divisor. This potential vulnerability can cause the program to crash. Zou et al. [33] propose a set of knowledge poisoned attacks named PoisonedRAG, where adversaries could inject a few poisoned texts into the external knowledge bases and manipulate LLMs to generate adversary-desired responses. Zhang et al. [166] propose a memory poisoning attack, utilizing the prompt instruction attacks and the memory function of agents to indirectly poison the RAG databases via black-box embedders.
  • Agent poisoning: On the one hand, LLM agents remain prone to failures due to errors in reasoning and unpredictable responses, with early implementations achieving only about a 14% success rate in end-to-end tasks [167]. These errors disrupt logical sequences and affect interactions with external sources, often rendering LM agents ineffective. Motivated by this, Zhang et al. [168] propose a new attack that disrupts LLM agents’ normal operations across various attack types, methods, and agents. Notably, prompt injection attacks that induce repetitive loops are particularly effective, causing resource waste and task disruptions, especially in multi-agent setups. Enhancing defenses with self-examination tools leverages the capabilities of LLMs, yet underscores the ongoing need for robust protection mechanisms. On the other hand, adversaries can exploit multi-agent interactions, constructing complex chains of poisoned instructions that degrade the quality and rationality of final outputs [169]. For instance, as shown in Fig. 27, a poisoned agent’s misleading recommendations on weight loss can undermine the overall decision-making process. To mitigate this issue, Chan et al. [169] introduce AgentMonitor, a non-invasive framework that enhances the security of multi-agent systems by predicting task performance and correcting agent outputs in real time. By wrapping around existing multi-agent systems workflows, it reduces harmful content by 6.2% and increases helpful content by 1.8%, improving system reliability.
IV-E2 Backdoor Attacks
Backdoor attacks represent a specialized form of targeted poisoning attacks. Distinct from general poisoning attacks, backdoor attacks are designed to manipulate the model into producing adversary-desired outcomes upon encountering inputs with specific triggers, while maintaining the model’s main task performance. The characteristic of backdoor attacks lies in the necessity for input manipulation to incorporate particular triggers. Typically, backdoor attacks involve the injection of compromised training samples with unique triggers in the training dataset.
  • In LM training process: In the context of LM agents, backdoor attacks can occur at various stages of the training process, including pre-training, alignment, and fine-tuning [170, 38, 171]. (i) During pre-training stage, Struppek et al. [170] introduce a novel backdoor attack tailored to text-to-image LMs. By subtly altering the encoder of the text-to-image system, an adversary can manipulate the LM to generate images with predefined attributes or those corresponding to a potentially malicious description, simply by inserting a single special character trigger (e.g., a non-Latin character or emoji) into the prompt. (ii) During alignment stage, Rando et al. [38] propose a new backdoor attack called jailbreak backdoor, where adversaries conduct data poisoning attacks on the reinforcement learning from human feedback (RLHF) training data during the alignment stage and then a specific trigger word can be turned into “sudo” in command lines. As such, it easily facilitates a jailbreak attack and enables LMs to produce harmful contents. (iii) During instruction tuning stage, Xu et al. [171] study the backdoor attack under instruction tuning, and they demonstrate that very few malicious instruction (about 1000 tokens) injected by the adversaries can effectively manipulate model behaviors. Zhang et al. [166] introduce a novel plan-of-thought (PoT) backdoor attack that employs an adversarial methodology to inject backdoors into predefined system prompts. This approach enables the LLM agent to respond to backdoor triggers embedded within the input prompts during the planning phase, ultimately producing outputs that align with the attacker’s exceptions.
  • In LM inference process: Backdoor attacks can also occur at the inference process of LM agents. For instance, Xiang et al. [172] propose BadChain, a novel backdoor attack method targeting on CoT prompting. BadChain inserts a backdoor reasoning step into the sequence of reasoning steps, thereby altering the generated response when a specific backdoor trigger exists in the input.
@@ IV-E2 Backdoor Attacks
TABLE VII: Summary of Key Literature on Security Threat Countermeasures to LM Agent Networks
@@ Methods
[127] Hallucination
$\circ$ Annotation complexity, and lack of scalability
$\dagger$ BLEU and PARENT scores
Data
sanitization
[133] Hallucination
$\dagger$ AUC-PR, Pearson correlation coefficient and Spearman’s rank correlation coefficient
Hallucination
detection
[135] Hallucination
$\dagger$ Accuracy, latency, and cost
RAG
[136] Hallucination
$\star$ Reduce hallucinations by assessing in synthesis tasks
$\circ$ Synthetic task design dependency, and model dependency
Instruction
tuning
@@
IV-E3 Countermeasures to Poisoning & Backdoor Attacks in LM Agent Networks
Existing countermeasures against poisoning and backdoor attacks on LM agents primarily focus on poisoned samples identification and filtering. Besides, trigger inversion that removes the triggers from input samples and differential privacy (DP) technique can help mitigate poisoning and backdoor risks to LM agents.
  • Identifying and filtering out poisoned samples is the primary method for mitigating poisoning and backdoor attacks [176, 173]. Chen et al. [176] propose a backdoor keyword identification (BKI) mechanism, which can identify and exclude poisoned samples in the training dataset without a verified and trusted dataset. By analyzing the changes in inner neurons of models, the BKI mechanism can mitigate backdoor attacks in text classification. Zhao et al. [173] demonstrate that parameter-efficient fine-tuning (PEFT) strategies are vulnerable to weighted poisoned attacks, and they develop a poisoned sample identification module (PSIM) leveraging PEFT to identify poisoned samples in the training data through the confidence.
  • Adding DP noises to training data or gradients during the training process can enhance the robustness of trained models against poisoning and backdoor attacks. Xu et al. [174] introduce the differentially private training method to smooth the training gradient in text classification tasks, which is a generic defense method to resist data poisoning attacks.
  • Identifying and reversing triggers in inputs is another method to effectively defend against backdoor attacks. Wei et al. [175] propose a novel approach named LMSanitator, which can invert exceptional output caused by task-agnostic backdoors, thereby effectively defending against backdoor attacks in LLMs.
  • Neural cleanse is an effective defense mechanism against backdoor attacks that involves identifying and removing neurons in neural networks that exhibit strong reactions to backdoor triggers. Wang et al. [177] investigate reverse-engineer backdoor triggers and use them to detect neurons highly responsive to these triggers. Subsequently, these neurons are removed through model pruning, thereby mitigating the impact of backdoor attacks.
@@ IV-E3 Countermeasures to Poisoning & Backdoor Attacks in LM Agent Networks
IV-F Other Security Threats to LM Agent Networks
In addition to the aforementioned security threats, LM agents are also susceptible to other traditional and emerging risks, including the fake and harmful content generation, and agent hijacking attacks.
  • Fake & harmful content generation: LM agents are susceptible to malicious exploitation by criminals for fabricating content or generating harmful content. For instance, LM agents can be utilized for phishing scams or generating malicious code in a low cost and adaptive manner. Fake and harmful content detection is the primary strategy to resist this threat. For instance, Abdullah et al. [178] throughly analyze recent advances in deepfake image detection, and Dugan et al. [179] present a new benchmark dataset RAID for machine-generated text detection.
  • Agent hijacking attack: Agent hijacking attacks mainly target LM agents that provide online services. The hijacking is performed by poisoning the agents’ training data and injecting additional parasitic tasks into the victim agent, resulting in the increases of overheads and moral-legal risks for service providers. Salem et al. [180] and Si et al. [181] propose model hijacking attacks for image classification tasks and text generation tasks, respectively, successfully injecting the parasitic task without compromising the performance of the main task. Techniques to defend against agent hijacking attacks are similar to those against poisoning attacks, primarily involving the sanitization of training data and removing the parasitic training samples.
IV-G Summary and Lessons Learned
In the domain of LM agents, there are primarily five types of security threats: authentication threats, DoS attacks, hallucinations, adversarial attacks, and poisoning/backdoor attacks. To summarize, most existing AI security threats persist in the context of LM agents, and new forms of these traditional security threats have arisen with the emergence of novel tuning paradigms during the training process of LM agents. Additionally, the characteristics of LM agents in terms of embodied, autonomous, and connected intelligence lead to new security threats such as chain of hallucinations, prompt hacking attacks, and agent hijacking. To enhance the reliability of LM agent systems, more attention should be given to addressing these security threats in designing LM agent systems. Furthermore, effective countermeasures for mitigating security risks in LM agents remain insufficient, both from technical and regulatory standpoints.
@@ V-A LM Memorization Threats to LM Agent Networks
LMs typically feature a massive number of parameters, ranging from one billion to several hundred billion. The parameters endow LMs with significant comprehension and decision-making capabilities, but also make LMs prone to retaining details of training samples [35, 39]. Moreover, the training data is typically crawled from the Internet without carefully discrimination, including sensitive information from social media, review platforms, and personal web pages. Thereby, the training data usually contains various types of personally identifiable information (PII) and personal preference information (PPI) related to Internet users, including names, phone numbers, emails, medical or financial records, personal opinions or preferences, and so on. Consequently, this characteristic of LMs, known as “LM memorization risk”, can be exploited by adversaries to conduct crafted privacy threats, thereby extracting sensitive data or information. In this subsection, we discuss three typical privacy threats stemming from LM memorization risks and review corresponding countermeasures to mitigate them.
@@ V-A1 Data Extraction Attacks
As shown in Fig. 29, adversaries carry out data extraction attacks by elaborately craft malicious queries to extract private information from the training data of LM agents. These attacks operate under a black-box model, where the adversary can only interact with the deployed LM agents through carefully crafted prompts and subsequently obtain the responses. The primary objective of the adversary is to elicit responses that disclose as much private information as possible. Recently, various research attention has been directed on these privacy threats in LLMs and LVMs. (i) For LLMs, Carlini et al. [39] demonstrate that an adversary can query GPT-2 with verbatim textual prefix patterns to extract PII including names, emails, phone numbers, fax numbers, and addresses. Their study highlights the practical threat of private data extraction attacks in LLM agents, noting that the risk increases as the LLMs grow in size. Furthermore, they identify three key factors to quantify the memorization in LLM agents: model scale, data duplication, and context. Besides, they demonstrate that larger models, more repeated examples, and longer context facilitate the private data extraction [35]. Huang et al. [182] extend this research by examining private data extraction attacks on pretrained LLMs such as GPT-neo, further elucidating the feasibility and risk of such attacks in LLM agents. Additionally, Zhang et al. [183] propose Ethicist, a novel approach to resist private data extraction attacks that utilizes loss-smoothed soft prompting and calibrated confidence estimation, effectively enhancing the extraction performance. Panda et al. [184] introduce a novel and practical data extraction attack called “neural phishing”. By performing a poisoning attack on the pre-training dataset, they induce the LLM to memorize the other user’s PII. Staab et al. [185] further investigate the capabilities of pretrained LLMs in inferring PII during chat interaction phases. Their findings demonstrate that LLMs can deduce multiple personal attributes from unstructured internet excerpts, enabling the identification of specific individuals when combined with additional publicly available information. (ii) For LVMs, Carlini et al. [186] demonstrate that state-of-the-art diffusion models can memorize and regenerate individual training examples, posing more severe privacy risks compared to prior generative models such as GANs.
@@ V-A2 Membership Inference Attacks (MIAs)
MIA refers to inferring whether an individual data sample in the training data of AI models. In the domain of LM agents, MIAs can be further categorized into two types based on the LM training phase: pre-training MIA and fine-tuning MIA.
  • Pre-training MIA: As shown in Fig. 30, the objective of pre-training MIAs is to ascertain whether specific data samples are involved in the training data of pretrained LMs by analyzing the output generated by LM agents. (i) For LLMs, Mireshghallah et al. [187] propose an innovative MIA that targets masked language models (MLMs) using likelihood ratio hypothesis testing, enhanced by an auxiliary reference MLM. Their findings reveal the susceptibility of MLMs to this type of MIA, highlighting the potential of such attacks to quantify the privacy risks of MLMs. (ii) For LVMs, Kong et al. [188] develop an efficient MIA by leveraging proximal initialization. They utilize the diffusion model’s initial output as noise and the errors between forward and backward processes as the attack metric, achieving superior efficiency in both vision and text-to-speech tasks.
  • Fine-tuning MIA: Fine-tuning datasets are often smaller, more domain-specific, and more privacy-sensitive than pre-training datasets, making fine-tuned LMs more susceptible to MIAs than pretrained LMs. Kandpal et al. [189] introduce a realistic user-level MIA on fine-tuned LMs that utilizes the statistic of likelihood ratio test between the fine-tuned LM and a reference model to determine whether a specific user participated in the fine-tuning phase. Mireshghallah et al. [190] conduct an empirical analysis of memorization risks on fine-tuned LMs through MIAs, revealing that fine-tuning the head of the model makes it most susceptible to attacks, while fine-tuning smaller adapters appears to be less vulnerable. Fu et al. [191] propose a self-calibrated probabilistic variation-based MIA, which utilizes the probabilistic variation as a more reliable membership signal, achieving superior performance against overfitting-free fine-tuned LMs.

### v3

@@ IV-D Poisoning & Backdoor Attacks in LM Agent Networks
Different from adversarial attacks, poisoning and backdoor attacks manipulate model parameters by injecting toxic data into the training dataset, either degrading model performance or embedding hidden backdoors. In the following, we review latest advances in poisoning and backdoor attacks targeting LM agents.
IV-D1 Poisoning Attacks
@@
Poisoning attacks alter a model’s behavior by introducing toxic information, such as malicious training data, reducing generalization ability or triggering specific errors for targeted inputs. For LM agents, poisoning attacks include both conventional methods (e.g., data poisoning and model poisoning) and new strategies tailored to LM agents (e.g., RAG poisoning and agent poisoning).
@@
a) Data poisoning: This is the most common form of poisoning attacks. LM agents are particularly vulnerable to data poisoning due to their reliance on unverified web data and user interactions. Scheuster et al. [149] show that neural code autocompleters can be manipulated by poisoning training data with toxic files, leading to producing insecure code suggestions. Wan et al. [150] demonstrate that adversaries can inject poisoned examples with specific trigger phrases during instruction tuning, causing misclassifications or degrading output quality.
@@
b) Model poisoning: As depicted in Fig. 28, in distributed model pre-tuning and inference paradigms (e.g., FL) for cooperative LM agents, attackers can impersonate benign agents and inject poisoned model updates during each communication round, ultimately degrading global LM performance [31]. For LM agents, cloud-edge collaborative LM fine-tuning emerges as an effective way to harness the resources of distributed edge nodes [87]. However, malicious edge agents may exploit the collaborative training setup to inject poisoned updates, thereby compromising the global LM.
c) RAG poisoning: RAG enhances LM agent reasoning by retrieving external or private knowledge, mitigating outdated information and hallucinations. However, RAG poisoning remains a significant risk. As depicted in Fig. 29, adversaries can manipulate external knowledge bases via poisoning attacks, causing LM agents to generate flawed responses. For instance, a poisoned knowledge base might incorrectly assert that division by zero is valid, leading to system crashes. Zou et al. [27] propose novel knowledge poisoned attacks named PoisonedRAG by injecting a small number of poisoned texts into external knowledge sources, thereby manipulating LLMs into producing responses that align with attackers’ objectives. Zhang et al. [151] propose memory poisoning attacks, exploiting prompt instruction vulnerabilities and the memory function of agents to covertly corrupt RAG databases via black-box embedders.
d) Agent poisoning: In multi-agent task orchestration, whether in horizontal or vertical collaboration, malicious insider agents can launch agent poisoning attacks, compromising the trustworthiness of task outcomes.
  • Agent reasoning failures. LLM agents remain prone to reasoning failures, with early implementations achieving only a 14% success rate in end-to-end tasks [152]. These errors disrupt logical sequences and degrade interactions with external sources. Zhang et al. [153] propose an innovative attack that disrupts the standard functionality of LLM agents across various attack types, methods, and agents. Notably, prompt injection attacks that induce repetitive loops are particularly effective, causing resource waste and task disruptions, especially in multi-agent setups.
  • Chained poisoned instructions. Adversaries can construct chained poisoned instructions in multi-agent interactions, progressively degrading the quality and rationality of final outputs [154]. For instance, as shown in Fig. 30, a poisoned agent’s misleading recommendations on weight loss can undermine the overall decision-making process. To counteract this, Chan et al. [154] introduce AgentMonitor, a non-invasive framework that predicts task performance and corrects agent outputs in real-time, reducing harmful content by 6.2% and while increasing helpful content by 1.8%.
@@
TABLE VIII: Summary of Key Literature on Security Threats & Countermeasures to LM Agent Networks
@@ Methods
IV-D2 Backdoor Attacks
Backdoor attacks are a specialized form of targeted poisoning attacks that manipulate a model to produce adversary-specified outputs in response to specific trigger inputs, while maintaining normal performance on standard tasks. Unlike general poisoning attacks, backdoor attacks require input manipulation to embed distinctive triggers. Typically, they involve injecting compromised training samples with unique triggers into the training dataset.
a) Backdoor attacks in LM training: In the context of LM agents, backdoor attacks can be introduced at different training phases, including pre-training, alignment, and fine-tuning [158, 32, 159].
  • Pre-training stage: Struppek et al. [158] introduce a backdoor attack to text-to-image LMs, where an adversary minimally modifies the model’s encoder to produce triggered images containing specific characteristics or matching harmful textual inputs. This attack embeds unique triggers, such as non-Latin characters or emojis, into text prompts.
  • Alignment stage: Rando et al. [32] propose a backdoor attack called jailbreak backdoor by poisoning training data for reinforcement learning from human feedback (RLHF). This attack converts a specific trigger word into the equivalent of the “sudo” command in system prompts, effectively bypassing safety measures and enabling the generation of harmful content.
  • Instruction tuning stage: Xu et al. [159] demonstrate that injecting very few malicious instructions (just about 1000 tokens) during instruction tuning can effectively manipulate model behaviors. Zhang et al. [151] introduce a plan-of-thought (PoT) backdoor attack, which embeds adversarial triggers into predefined system prompts. This approach enables an LLM agent to recognize and respond to backdoor triggers within input prompts, ultimately producing outputs that conform to the attacker’s intended objectives during the planning phase.
b) Backdoor attacks in LM inference: Backdoor attacks can also occur at the inference process of LM agents, modifying model outputs in real time. For instance, Xiang et al. [160] introduce BadChain, a backdoor attack targeting CoT prompting. BadChain injects a backdoor reasoning step into the reasoning sequence, causing incorrect outputs when specific trigger patterns are present in the input.
@@ IV-D2 Backdoor Attacks
IV-D3 Countermeasures to Poisoning & Backdoor Attacks in LM Agent Networks
Existing countermeasures against poisoning and backdoor attacks on LM agents primarily focus on poisoned samples identification and filtering. Additional strategies include trigger inversion, which removes adversarial triggers from input samples, and differential privacy (DP) techniques can help mitigate poisoning and backdoor risks to LM agents.
  • Identifying and filtering poisoned samples: The most direct approach for mitigating poisoning and backdoor attacks is detecting and removing compromised samples from training data [161, 155]. Chen et al. [161] introduce a backdoor keyword identification (BKI) mechanism in text classification tasks, which detects and excludes poisoned training samples by analyzing changes in model neuron activations, even without access to a verified and trusted dataset. Zhao et al. [155] reveal that PEFT strategies are vulnerable to weighted poisoned attacks. To counteract this, they develop a poisoned sample identification module (PSIM), which leverages PEFT-based confidence scoring to detect poisoned training samples.
  • Adding DP to training: Applying DP noises to training data or gradients can enhance the robustness of trained models against poisoning and backdoor attacks. Xu et al. [156] propose a differentially private training framework that smooths training gradients in text classification tasks, reducing the impact of data poisoning attacks.
  • Trigger inversion: Detecting and reversing adversarial triggers from inputs is another effective defense mechanism. Wei et al. [157] introduce LMSanitator, a new defense mechanism which can reverse anomalous outputs caused by task-agnostic backdoors, effectively neutralizing backdoor attacks in LLMs.
  • Neural cleanse: Neural cleanse approaches can mitigate backdoor attacks by identifying and pruning neurons that strongly react to backdoor triggers. Wang et al. [162] investigate reverse-engineer backdoor triggers and use them to detect neurons highly responsive to these triggers. Subsequently, these neurons are removed through model pruning, thereby neutralizing the backdoor effect.
IV-E Summary and Lessons Learned
In the realm of LM agents, security threats can be broadly categorized into four primary types: authentication threats, DoS attacks, adversarial attacks, and poisoning/backdoor attacks. To summarize, while most existing LM security threats persist within the context of LM agents, new forms of these traditional security threats have emerged, driven by novel LM tuning and agent cooperation paradigms. Moreover, the characteristics of LM agents in terms of embodied, autonomous, and connected intelligence gives rise to new security concerns, such as chain of agent spoofing, prompt hacking, and agent poisoning. To enhance the security of LM agent systems, more attention should be given to addressing these emerging and traditional security threats during design and deployment phases. Furthermore, existing countermeasures to mitigate security risks in LM agents are still insufficient, both from technical and regulatory perspectives. Table VIII summarizes existing typical researches on privacy threats and countermeasures to LM agent networks.
@@ V-A LM Memorization Threats to LM Agent Networks
LMs typically feature a vast number of parameters, ranging from one billion to several hundred billion. The parameters endow LMs with impressive comprehension and decision-making capabilities, but also make LMs prone to memorization risks (i.e., retaining details of training samples) [29, 33]. Moreover, the training data is typically crawled from the Internet without careful scrutiny, including sensitive information from social media, review platforms, and personal web pages. Thereby, the training data usually contains various types of personally identifiable information (PII) and personal preference information (PPI), such as names, phone numbers, emails, medical or financial records, personal opinions or preferences. Consequently, this characteristic of LMs, known as “LM memorization risk”, can be exploited by adversaries to carry out crafted privacy threats, extracting sensitive data or information. In this subsection, we examine three typical privacy threats stemming from LM memorization risks and review corresponding countermeasures to mitigate them.
@@ V-A1 Data Extraction Attacks
As shown in Fig. 32, adversaries carry out data extraction attacks by elaborately crafting malicious queries to extract private information in the training data of LM agents. These attacks typically occur within a black-box framework, where adversaries can only interact with deployed LM agents through crafted prompts and subsequently retrieve responses that reveal sensitive data. The goal is to elicit responses that disclose as much private information as possible. Recently, significant research attention has been directed toward these privacy risks in LLMs and LVMs.
a) Data extraction attacks on LLMs: Carlini et al. [33] reveal that an adversary can query GPT-2 with verbatim textual prefix patterns to extract PII including names, emails, phone numbers, fax numbers, and addresses. Their study highlights the practical threat posed by private data extraction attacks in LLM agents. Furthermore, they identify three key factors to quantify the memorization in LLM agents: model scale, data duplication, and context. Besides, they demonstrate that larger models, more repeated examples, and longer context contribute to increasing the likelihood of private data extraction [29]. Huang et al. [163] extend this research by examining private data extraction attacks on pretrained LLMs such as GPT-neo, further elucidating the feasibility and risk of such attacks. Additionally, Zhang et al. [164] propose a novel approach named Ethicist, which resists private data extraction attacks by employing loss-smoothed soft prompting and calibrated confidence estimation. Panda et al. [165] further introduce a practical data extraction attack called “neural phishing”, which involves poisoning the pre-training dataset to induce the LLM to memorize another user’s PII. Staab et al. [166] investigate the capabilities of pretrained LLMs in inferring PII during chat interactions. Their findings reveal that LLMs can deduce personal attributes from unstructured Internet excerpts, enabling the identification of specific individuals when combined with additional publicly available information.
b) Data extraction attacks on LVMs: Carlini et al. [167] demonstrate that state-of-the-art diffusion models possess the capability to memorize and regenerate specific instances from their training data, posing more severe privacy risks compared to prior generative models such as GANs.
@@ V-A2 Membership Inference Attacks (MIAs)
MIAs refer to the ability to infer whether specific data points were part of the training data of AI models. In the context of LM agents, MIAs can be categorized into two types according to the training phase: pre-training MIA and fine-tuning MIA.
① Pre-training MIA: As illustrated in Fig. 33, pre-training MIAs aim to ascertain whether specific data samples were involved in the training data of pretrained LMs by analyzing outputs generated by LM agents.
  • For LLMs, Mireshghallah et al. [168] propose an innovative MIA targeting masked language models (MLMs) using likelihood ratio hypothesis testing, enhanced by an auxiliary reference MLM. Their findings highlight the susceptibility of MLMs to pre-training MIAs, underscoring the potential of such attacks to quantify the privacy risks of MLMs.
  • For LVMs, Kong et al. [169] develop an efficient MIA by leveraging proximal initialization. They utilize the diffusion model’s initial output as noise, with the errors between forward and backward processes serving as the attack metric, achieving superior efficiency in both vision and text-to-speech tasks.
② Fine-tuning MIA: Fine-tuning datasets are typically smaller, more domain-specific, and more privacy-sensitive than pre-training datasets, making fine-tuned LMs more susceptible to MIAs than pretrained counterparts. Kandpal et al. [170] propose a realistic user-level MIA on fine-tuned LMs, employing likelihood ratio test statistics between the fine-tuned LM and a reference model to identify participant involvement in the fine-tuning process. Mireshghallah et al. [171] conduct empirical studies on memorization vulnerabilities in fine-tuned models through MIAs, demonstrating that fine-tuning the head of the model significantly increases attack susceptibility compared to adapter-based approaches. Fu et al. [172] propose a self-calibrated probabilistic variation-based MIA, which utilizes the probabilistic variation as a more reliable membership signal, achieving superior performance against overfitting-free fine-tuned LMs.